# Anosov diffeomorphisms of open surfaces

### v1

@@ 4.3. Proper Global Leaves
In this subsection we define a family of global unstable leaves with many useful properties with respect to the coding, the margulis measures, their density in $M$, and such that this family of global unstable leaves covers the residual set $Y$. These properties will later be used to extend the Margulis measures in a unique way that makes them holonomy invariant across different elements of the Markov partition.
@@
A chain $\underline{R}\in\widehat{\Sigma}_{L}$ is called standard if there exists $x\in\widehat{\pi}[\tau^{-1}[\{\underline{R}\}]]$ such that $W^{u}(x,R_{0})$ is a standard unstable leaf (recall Definition 3.16).
@@
For any $\underline{R}^{-}\in\widehat{\Sigma}_{L}^{\circ}$ and any $\underline{R}^{+}\in\widehat{\Sigma}_{R}$ with $R^{+}_{0}=R^{-}_{0}$, the word $\underline{R}^{-}\underline{R}^{+}$ is in $\widehat{\Sigma}^{\circ}$, and hence the point coded by this word is in $Y$.
@@
Let $\mathcal{C}_{u}$ denote the the curve $W^{u}(\underline{R}^{-})$ in the Markov rectangle $R_{0}$. The Local product coordinates on $R_{0}$ make $W^{u}(\underline{R}^{-})$ look like a horizontal line in the box $[0,1]\times[0,1]$ while the curve $\mathcal{C}_{s}$ corresponding to $W^{s}(\underline{R}^{+})$ looks like a vertical line. Note that the assumptions on $\underline{R}^{-}$ and $\underline{R}^{+}$ imply that these vertical and horizontal lines do not coincide with the boundary of the square $[0,1]\times[0,1]$; thus we see that the intersection of $W^{u}(\underline{R}^{-})$ and $W^{s}(\underline{R}^{+})$ is a point in the interior of $R_{0}$, and the result follows. ∎
@@
As there are at most countably many nonstandard chains, it suffices to show that if $\underline{R}$ is a standard chain, then $\underline{R}\in\widehat{\Sigma}^{\circ}_{L}$. As a reminder, a chain $\underline{R}$ is standard if $W^{u}(\underline{R})$ does not intersect the orbit of the unstable boundaries of the Markov rectangles. Suppose that $\underline{R}$ is such a chain and let $\underline{Q}\in\tau^{-1}[\{\underline{R}\}]$ and $x=\widehat{\pi}(\underline{Q})$. Then there is a point $y\in W^{s}(x,R)$ such that $W^{s}(y)$ does not meet the orbit of the stable boundaries. Hence we see that $[x,y]$ lies in $Y^{\prime}$. Note that if $\underline{S}$ is a preimage of $y$ under $\widehat{\pi}$, then as $\widehat{\pi}([\underline{Q},\underline{S}])=[\widehat{\pi}(\underline{Q}),\widehat{\pi}(\underline{S})]$. As $z\in Y^{\prime}$, we see that $[\underline{Q},\underline{S}]\in\widehat{\Sigma}^{\circ}$ and so $[\underline{Q},\underline{S}]\in\widehat{\Sigma}^{\circ}_{L}$. ∎
@@
For $\underline{R}\in\widehat{\Sigma}$, let $W^{u}(\underline{R})$ denote the local unstable leaf $W^{u}(x,R_{0})$ where $x$ is any point in $\hat{\pi}(\tau^{-1}(\{\underline{R}\}))$.
@@
A global unstable leaf $W^{u}$ in $M$ is called proper if there exists a proper chain $\underline{R}\in\widehat{\Sigma}_{L}^{\star}$ and $x\in\hat{\pi}(\tau^{-1}\{\underline{R}\})$ such that $W^{u}$ contains $x$.
@@ 4.4. Global Holonomy Invariance of the Margulis Measures
So far we have constructed a family of measures on proper global leaves, which are dense in the manifold. In order to extend this family of measures coherently to the whole manifold, we need to establish global continuity of these measures under stable holonomy. That is, we need to establish continuity on the manifold for holonomies that do lie within a single Markov rectangle. In order to show this, we will do what amounts to a specific calculation which is enabled by use of the explicit combinatorial formula for the harmonic function from §4.1. The important feature of these harmonic functions is that they are defined, in a sense, dynamically because they can be interpreted as counting the number of intersections between stable and unstable manifolds. The importance of the holonomy invariance of the measure will become evident later, when we use this property to construct the universal cover of $M$.
This subsection contains two proofs, one for the transient case and one for the recurrent case. We begin with the transient case. Note that we only prove holonomy invariance for proper leaves. Before we give the proof, we give a brief sketch of the ideas it involves.
@@
The idea of the proof is as follows. Recall the construction of the harmonic functions on the shift space $\widehat{\Sigma}_{L}$. These were constructed by counting the number of words of length $n$ between two symbols $R_{1},R_{2}\in\mathcal{R}$. Recall that $Z^{\prime}_{n}(R_{1},R_{2})$ counts precisely this quantity. The measure of a set $U\subset W^{u}(x,R)$ is then determined by looking at the preimage of this set, covering its preimage with cylinders whose measure is precisely determined by counting words passing through that cylinder. In order to make proper use of these measures, we must relate this purely combinatorial description with the topological facts of the dynamics in the manifold. The key observation is that if $W^{u}(x,R_{1})$ is standard then we can estimate the quantity $Z^{\prime}_{n}(R_{1},R_{2})$ by counting the intersections of $W^{u}(x,R_{1})$ with $f^{-n}(W^{s}(y,R_{2}))$. If we choose $x$ and $y$ correctly, then $\left|W^{u}(x,R_{1})\right|\cap f^{-n}(W^{s}(y,R_{2}))$ differ by at most $2$. Hence we can estimate the measure of subsets of $W^{u}(x,R)$ by first covering them with intervals and then counting their intersections with $f^{-n}(W^{s}(y,R_{2}))$. The important part of this is that if $R^{\prime}$ is a rectangle sitting so that stable leaves in the rectangle $R_{1}$ exit $R_{1}$ into $R^{\prime}$, then the if $I_{1}\subset W^{u}(x,R_{1})$ is an interval and $I_{2}\subset W^{u}(x,R^{\prime})$ are two holonomy related intervals, then the number of intersections between $f^{-n}(W^{s}(y,R_{2}))$ and $I_{1}$ should be almost equal to the number of intersections of $f^{-n}(W^{s}(y,R_{2}))$ with $I_{2}$. Consequently, one expects that the measures $\mu_{\underline{R}}$ to have holonomy invariance even when the holonomies pass between different rectangles. This is only possible because we have constructed the harmonic functions in terms of what is essentially dynamical information.
@@
Assume that $\widehat{\Sigma}_{L}$ is transient, then the Margulis measures on the proper global unstable leaves are holonomy invariant. Specifically, let $R,S\in\mathcal{R}$ be adjacent rectangles such that $L\coloneqq\partial^{u}R\cap\partial^{u}S\neq\varnothing$. Let $\underline{R}\in[R]\cap\widehat{\Sigma}_{L}^{\star},\underline{S}\in[S]\cap\widehat{\Sigma}_{L}^{\star}$, and let $x_{R}\in\widehat{\pi}[\tau^{-1}[\{\underline{R}\}]]$, $x_{S}\in\widehat{\pi}[\tau^{-1}[\{\underline{S}\}]]$. Let $V^{s}(L)$ be the saturation of local stable leaves in $R$ and $S$ of $L$, and set $I_{R}\coloneqq V^{s}(L)\cap W^{u}(x_{R},R)$ and $I_{S}:=V^{s}(L)\cap W^{u}(x_{S},S)$. Then $\mu_{W^{u}(\underline{R})}|_{I_{R}}=\mu_{W^{u}(\underline{S})}|_{I_{S}}\circ\Gamma$, where $\Gamma$ is the local holonomy along the stable foliation restricted to $V^{s}(L)$. The same holds in the case that $\widehat{\Sigma}_{L}$ is recurrent.
@@
If $x\in W^{u}(x_{R},R)\cap W^{s}(x_{k})$, then $x$ has a coding in $\tau^{-1}[\{\underline{R}\}]$ and this coding is unique.
@@
We begin by showing that a coding exists. To begin, take a sequence of points $y_{n}\in Y^{\prime}$ such that $y_{n}\to x$. Let $\underline{R}^{(n)}$ denote the coding of these points, which lies in $\widehat{\Sigma}^{\circ}$. Then consider the points $W_{n}=(R_{i})_{i\leq 0}\cdot(R_{i}^{(n)})_{i\geq 0}$, which lie in $\widehat{\Sigma}^{\circ}$ by Claim 4.21. By the local compactness of $\widehat{\Sigma}$, some subsequence converges to a point $W\in\tau^{-1}(\{\underline{R}\})$. Note then that this point codes $x$ by continuity as each point $W_{n}$ codes the point $[x_{R},y_{n}]\in W^{u}(x_{R},R)$.
We now show that this coding is unique. Let $\underline{u}^{1}$ and $\underline{u}^{2}$ be two codings in $\tau^{-1}(\{\underline{R}\})$ of a point $x\in W^{u}(x_{R},R)\cap W^{s}(x_{k})$. Note that sufficiently far in the future $\underline{u}^{1}$ and $\underline{u}^{2}$ are equal as they approach $x_{k}$, whose orbit always lies in the interior of Markov rectangles. As these codings both lie in $\tau^{-1}(\{\underline{R}\})$, the agree in the past as well. Hence by Lemma 3.19 $\underline{u}^{1}=\underline{u}^{2}$. ∎
We will now show the holonomy invariance of $\mu_{W^{u}(\underline{R})}$. We begin with what is the most difficult case, an atom that lies on the common boundary of $R$ and $T$. This argument contains all the difficulties that occur in the remaining cases. We begin by finding a formula for a measure of a cylinder in terms of counting intersections of a stable and unstable manifold.
@@
Suppose that $\underline{R}\in\widehat{\Sigma}_{L}$ and that that $C$ is a cylinder contained in $\tau^{-1}(\{\underline{R}\})$. Then
$\left|Z_{i}(\underline{C},\omega_{n_{k}})\right|=\left|\widehat{\pi}(C))\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))\right|.$
@@
Each point $x$ in $\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))$ is uniquely coded by Claim 4.28. Further, the map that carries $x$ to its coding in $\tau^{-1}(\{\underline{R}\})$ is a surjection. This follows because by the Markov property any word $w\in Z_{i}(\underline{C},\omega_{n_{k}})$ may be extended to a two sided word $\underline{w}\in\widehat{\Sigma}$ lying in the cylinder $[\underline{c}]$ such that $\underline{w}$’s $n$th iterate lies in the partition set $[\omega_{n_{k}}]$ in the local stable manifold of the periodic point $x_{k}$. Hence the image of this point lies in $\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))$. ∎
An important and useful property that we will use is that the image of a cylinder $[C]\subset\tau^{-1}(\{\underline{R}\})$ is a closed interval.
@@
We have the following formula for the measure of a cylinder $C=[C_{1},\ldots,C_{N}]$,
$\widehat{\mu}_{\underline{S}}([c_{1},\ldots,c_{n}])=\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
We first will obtain a combinatorial forumla for the measure of a cylinder set.
By (2) in Corollary 4.11, which gives for a cylinder of length $1$ that:
$\widehat{\mu}_{\underline{S}}([c])=\mu_{\underline{S}}(1_{[\underline{C}]})\circ\sigma)=e^{-h_{G}(\sigma)}\sum_{\sigma_{R}\underline{S}=\underline{R}}\widehat{\mu}_{\underline{S}\underline{C}}(1_{[C_{-1}]})=e^{-h_{G}(\sigma)}\widehat{\mu}_{\underline{S}\underline{c}}(1),$
where the last equality follows because only one term in the sum assigns any measure to the integrand. Hence for a cylinder $[\underline{c}]$ of length $N$,
$\widehat{\mu}_{\underline{S}}([\underline{c}])=e^{-Nh_{G}(\sigma)}\widehat{\mu}_{\underline{S}\underline{c}}(1).$
Then, by definition, we can rewrite this as
$\widehat{\mu}_{\underline{S}}([\underline{c}])$ $=e^{-Nh_{G}(\sigma)}\lim_{k}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(\underline{C_{N}},\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z(\underline{a},\omega_{n_{k}})}$
$=e^{-Nh_{G}(\sigma)}\lim_{k}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i+N}^{\prime}(\underline{C},\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z(\underline{a},\omega_{n_{k}})}$
$=\lim_{k}\frac{O(1)+\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(\underline{C},\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z(\underline{a},\omega_{n_{k}})},$
where the $O(1)$ is uniformly bounded because as $k\to\infty$, arbitrarily many of the finite number of terms that were gained when the sum was reindexed are zero as $\omega_{n_{k}}$ moves away from $\widehat{\pi}(\underline{C})$ in the manifold. We now apply Claim 4.29 to conclude. ∎
@@
$\mathcal{C}^{N}_{R}(U)=\{\text{all cylinders }[\underline{c}]\text{ in }\tau^{-1}(\{\underline{R}\})\text{ of length }N\text{ such that }[\underline{c}]\cap\widehat{\pi}^{-1}[\{x\}]\neq\emptyset\}$
In addition we define the set
@@
$\mu_{\underline{R}}(\{x\})=$ $\inf_{N}\sum_{[C]\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}([C]).$
@@
$=\sum_{C\in\mathcal{C}_{N}^{S}(\{\Gamma(x)\})}\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(\mathcal{D}_{N}^{S})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$

### v2

@@ 3.4. Proper Global Leaves
In this subsection we define a family of global unstable leaves with many useful properties with respect to the coding, the Margulis measures, their density in $M$, and such that this family of global unstable leaves covers the residual set $Y$’ from Definition 2.14. These properties will later be used to extend the Margulis measures in a unique way that makes them holonomy invariant across different elements of the Markov partition.
@@
A chain $\underline{R}\in\widehat{\Sigma}_{L}$ is called standard if there exists $x\in\widehat{\pi}[W^{u}_{0}(\underline{R})]$ such that $W^{u}(x,R_{0})$ is a standard unstable leaf (recall Definition 2.15).
@@
For any $\underline{R}^{-}\in\widehat{\Sigma}_{L}^{\circ}$ and any $\underline{R}^{+}\in\widehat{\Sigma}_{R}^{\circ}$ with $R^{+}_{0}=R^{-}_{0}$, the word $\underline{R}^{-}\underline{R}^{+}$ is in $\widehat{\Sigma}^{\circ}$, and hence the point coded by this word is in $Y$.
@@
Let $\mathcal{C}_{u}$ denote the portion of the curve $W^{u}(\underline{R}^{-})$ in the Markov rectangle $R_{0}$. The Local product coordinates on $R_{0}$ make $W^{u}(\underline{R}^{-})$ look like a horizontal line in the box $[0,1]\times[0,1]$ while the curve $\mathcal{C}_{s}$ corresponding to $W^{s}(\underline{R}^{+})$ looks like a vertical line. Note that the assumptions on $\underline{R}^{-}$ and $\underline{R}^{+}$ imply that these vertical and horizontal lines do not coincide with the boundary of the square $[0,1]\times[0,1]$; thus we see that the intersection of $W^{u}(\underline{R}^{-})$ and $W^{s}(\underline{R}^{+})$ is a point in the interior of $R_{0}$, and the result follows. ∎
@@
As there are at most countably many nonstandard chains, it suffices to show that if $\underline{R}$ is a standard chain, then $\underline{R}\in\widehat{\Sigma}^{\circ}_{L}$. As a reminder, a chain $\underline{R}$ is standard if $W^{u}(\underline{R})$ does not intersect the orbit of the unstable boundaries of the Markov rectangles. Suppose that $\underline{R}$ is such a chain and let $\underline{Q}\in\tau^{-1}[\{\underline{R}\}]$ and $x=\widehat{\pi}(\underline{Q})$. Then there is a point $y\in W^{s}(x,R)$ such that $W^{s}(y)$ does not meet the orbit of the stable boundaries. Hence we see that $[x,y]$ lies in the set $Y^{\prime}$ from Definition 2.14. Note that if $\underline{S}$ is a preimage of $y$ under $\widehat{\pi}$, then $\widehat{\pi}([\underline{Q},\underline{S}])=[\widehat{\pi}(\underline{Q}),\widehat{\pi}(\underline{S})]$. As $[x,y]\in Y^{\prime}$, we see that $[\underline{Q},\underline{S}]\in\widehat{\Sigma}^{\circ}$ and so $[\underline{Q},\underline{S}]\in\widehat{\Sigma}^{\circ}_{L}$. ∎
@@
A global unstable leaf $W^{u}$ in $M$ is called proper if there exists a proper chain $\underline{R}\in\widehat{\Sigma}_{L}^{\star}$ and $x\in\hat{\pi}(W^{u}_{0}(\underline{R}))$ such that $W^{u}$ contains $x$.
@@ 3.5. Global Holonomy Invariance of the Margulis Measures
So far we have constructed a family of measures on proper global leaves, which are dense in the manifold. In order to extend this family of measures coherently to the whole manifold, we need to establish global continuity of these measures under stable holonomy. That is, we need to establish continuity on the manifold for holonomies that do lie within a single Markov rectangle. In order to show this, we will do what amounts to a specific calculation which is enabled by use of the explicit combinatorial formula for the harmonic function from §3.2. The important feature of these harmonic functions is that they are defined, in a sense, dynamically because they can be interpreted as counting the number of intersections between stable and unstable manifolds. Note that we only prove holonomy invariance for proper leaves. Before we give the proof, we give a brief sketch of the ideas it involves.
@@
The idea of the proof is as follows. Recall the construction of the harmonic functions on the shift space $\widehat{\Sigma}_{L}$. These were constructed by counting the number of words of length $n$ between two symbols $R_{1},R_{2}\in\mathcal{R}$. Recall that $Z^{\prime}_{n}(R_{1},R_{2})$ counts precisely this quantity. The measure of a set $U\subset W^{u}(x,R)$ is then determined by looking at the preimage of this set and covering its preimage with cylinders whose measure is precisely determined by counting words passing through that cylinder. In order to make proper use of these measures, we must relate this purely combinatorial description with the topological facts of the dynamics in the manifold. The key observation is that if $W^{u}(x,R_{1})$ is standard then we can estimate the quantity $Z^{\prime}_{n}(R_{1},R_{2})$ by counting the intersections of $W^{u}(x,R_{1})$ with $f^{-n}(W^{s}(y,R_{2}))$. If we choose $x$ and $y$ correctly, then $\left|W^{u}(x,R_{1})\right|\cap f^{-n}(W^{s}(y,R_{2}))$ differ by at most $2$. Hence we can estimate the measure of subsets of $W^{u}(x,R)$ by first covering them with intervals and then counting their intersections with $f^{-n}(W^{s}(y,R_{2}))$. The important part of this is that if $R^{\prime}$ is a rectangle sitting so that stable leaves in the rectangle $R_{1}$ exit $R_{1}$ into $R^{\prime}$, then if $I_{1}\subset W^{u}(x,R_{1})$ is an interval and $I_{2}\subset W^{u}(x,R^{\prime})$ are two holonomy related intervals, then the number of intersections between $f^{-n}(W^{s}(y,R_{2}))$ and $I_{1}$ should be almost equal to the number of intersections of $f^{-n}(W^{s}(y,R_{2}))$ with $I_{2}$. Consequently, one expects that the measures $\mu_{\underline{R}}$ to have holonomy invariance even when the holonomies pass between different rectangles. This is only possible because we have constructed the harmonic functions in terms of what is essentially dynamical information.
We remark that [BM77] similarly push forward a measure from a symbolic system to a smooth system and then must check that this pushforward has holonomy invariance. A difference between their case and the present one is that the measures constructed in are non-atomic. Hence they are able to freely discard many troublesome points such as the boundaries of rectangles [BM77, Lemma 4.3]. On the other hand, we specifically consider the case that the Margulis measures have an atom that lies on the boundary of a rectangle. This is just one of the complications that appears in applying the strategy of Bowen-Marcus.
@@
Assume that $\widehat{\Sigma}_{L}$ is transient, then the Margulis measures on the proper global unstable leaves are holonomy invariant. Specifically, let $R,S\in\mathcal{R}$ be adjacent rectangles such that $L\coloneqq\partial^{u}R\cap\partial^{u}S\neq\varnothing$. Let $\underline{R}\in[R]\cap\widehat{\Sigma}_{L}^{\star},\underline{S}\in[S]\cap\widehat{\Sigma}_{L}^{\star}$, and let $x_{R}\in\widehat{\pi}[W^{u}_{0}(\underline{R})]$, $x_{S}\in\widehat{\pi}[W^{u}_{0}(\underline{S})]$. Let $V^{s}(L)$ be the saturation of local stable leaves in $R$ and $S$ of $L$, and set $I_{R}\coloneqq V^{s}(L)\cap W^{u}(x_{R},R)$ and $I_{S}:=V^{s}(L)\cap W^{u}(x_{S},S)$. Then $\mu_{W^{u}(\underline{R})}|_{I_{R}}=\mu_{W^{u}(\underline{S})}|_{I_{S}}\circ\Gamma$, where $\Gamma$ is the local holonomy along the stable foliation restricted to $V^{s}(L)$. The same holds in the case that $\widehat{\Sigma}_{L}$ is recurrent.
@@
If $x\in W^{u}(x_{R},R)\cap W^{s}(x_{k})$, then $x$ has a coding in $W^{u}_{0}(\underline{R})$ and this coding is unique.
@@
We begin by showing that a coding exists. To begin, take a sequence of points $y_{n}\in Y^{\prime}$ such that $y_{n}\to x$. Let $\underline{R}^{(n)}$ denote the coding of these points, which lies in $\widehat{\Sigma}^{\circ}$. Then consider the points $W_{n}=(R_{i})_{i\leq 0}\cdot(R_{i}^{(n)})_{i\geq 0}$, which lie in $\widehat{\Sigma}^{\circ}$ by Claim 3.20. By the local compactness of $\widehat{\Sigma}$, some subsequence converges to a point $W\in W^{u}_{0}(\underline{R})$. Note then that this point codes $x$ by continuity as each point $W_{n}$ codes the point $[x_{R},y_{n}]\in W^{u}(x_{R},R)$.
We now show that this coding is unique. Let $\underline{u}^{1}$ and $\underline{u}^{2}$ be two codings in $W^{u}_{0}(\underline{R})$ of a point $x\in W^{u}(x_{R},R)\cap W^{s}(x_{k})$. Note that sufficiently far in the future $\underline{u}^{1}$ and $\underline{u}^{2}$ are equal as they approach $x_{k}$, whose orbit always lies in the interior of Markov rectangles. As these codings both lie in $W^{u}_{0}(\underline{R})$, the agree in the past as well. Hence by Lemma 2.18 $\underline{u}^{1}=\underline{u}^{2}$. ∎
We will now show the holonomy invariance of $\mu_{W^{u}(\underline{R})}$. For simplicity, we will consider in detail only the most difficult case: when $\mu_{W^{u}(\underline{R})}$ has an atom that lies on the common stable boundary of $R$ and another rectangle $T$ that also abuts $S$. Note that we are specializing to a specific configuration of rectangles. This argument contains all the difficulties that occur in the remaining cases. We begin by finding a formula for a measure of a cylinder in terms of counting intersections of a stable and unstable manifold.
Previously, when we defined the harmonic function, we had the notation $Z_{i}^{\prime}(R_{1},R_{2})$ that counted the number of paths between $R_{1}$ and $R_{2}$ of length $i$. Now for a cylinder $C=[C_{1},\ldots,C_{N}]$ and $\omega_{n_{k}}\in\mathcal{R}$, we extend the definition of $Z_{i}^{\prime}$ so that $Z_{i}^{\prime}(C,\omega_{n_{k}})$ counts the number of paths from $C_{1}$ to $\omega_{n_{k}}$ of length $i$ of the form $C_{1},C_{2},\ldots,C_{N},\ldots,\omega_{n_{k}}$. Note that as long as $i$ is sufficiently large this is equal to the number of paths from $C_{N}$ to $\omega_{n_{k}}$ of length $i-N+1$.
@@
Suppose that $\underline{R}\in\widehat{\Sigma}_{L}$ and that that $C$ is a cylinder contained in $W^{u}_{0}(\underline{R})$. Then
$\left|Z_{i}^{\prime}(C,\omega_{n_{k}})\right|=\left|\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))\right|.$
@@
Each point $x$ in $\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))$ is uniquely coded by Claim 3.27. Further, the map that carries $x$ to its coding in $W^{u}_{0}(\underline{R})$ is a surjection. This follows because by the Markov property any word $w\in Z_{i}(C,\omega_{n_{k}})$ may be extended to a two sided word $\underline{w}\in\widehat{\Sigma}$ lying in the cylinder $C$ such that $\underline{w}$’s $n_{k}$th iterate lies in the partition set $[\omega_{n_{k}}]$ in the local stable manifold of the periodic point $x_{k}$. Hence the image of this point lies in $\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{n_{k}}))$. ∎
An important and useful property that we will use is that the image of a cylinder $C\subset W^{u}_{0}(\underline{S})$ is a closed interval.
@@
We have the following formula for the measure of a cylinder $C$,
$\widehat{\mu}_{\underline{S}}(C)=\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}_{0}(\underline{a}))\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
We first will obtain a combinatorial formula for the measure of a cylinder set.
By (2) in Corollary 3.10, which gives for a cylinder $[C_{0}]$ of length $1$ that:
$\widehat{\mu}_{\underline{S}}([C_{0}])=\widehat{\mu}_{\underline{S}}(1_{[C_{0}]}\circ\sigma)=e^{-h_{G}(\sigma)}\sum_{\sigma_{R}\underline{S}=\underline{R}}\widehat{\mu}_{\underline{SC_{0}}}(1_{[C_{0}]}\circ\sigma)=e^{-h_{G}(\sigma)}\widehat{\mu}_{\underline{SC_{0}}}(1),$
where the last equality follows because only one term in the sum assigns any measure to the integrand. Hence for a cylinder $C$ of length $N$,
$\widehat{\mu}_{\underline{S}}(C)=e^{-Nh_{G}(\sigma)}\widehat{\mu}_{\underline{SC}}(1).$
Then, by definition, as the chain is transient we can rewrite this as
$\widehat{\mu}_{\underline{S}}(C)$ $=e^{-Nh_{G}(\sigma)}\hat{\mu}_{\underline{SC}}(1)$
$=e^{-Nh_{G}(\sigma)}\lim_{k}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(C_{N},\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(\underline{a},\omega_{n_{k}})}$
$=e^{-Nh_{G}(\sigma)}\lim_{k}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i+N}^{\prime}(C,\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(\underline{a},\omega_{n_{k}})}$
$=\lim_{k}\frac{O(1)+\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(C,\omega_{n_{k}})}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}Z_{i}^{\prime}(\underline{a},\omega_{n_{k}})},$
where the $O(1)$ is uniformly bounded because as $k\to\infty$, arbitrarily many of the finite number of terms that were gained when the sum was reindexed are zero as $\omega_{n_{k}}$ moves away from $\widehat{\pi}(C)$ in the manifold. We now apply Claim 3.28 to conclude. ∎
@@
$\mathcal{C}^{N}_{R}(U)=\{\text{all cylinders }C\text{ in }W^{u}_{0}(\underline{R})\text{ of length }N\text{ such that }C\cap\widehat{\pi}^{-1}[\{x\}]\neq\emptyset\}.$
In addition, we define the set
@@
$\mu_{\underline{R}}(\{x\})=$ $\inf_{N}\sum_{C\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}(C).$
@@
$=\sum_{C\in\mathcal{C}_{N}^{S}(\{\Gamma(x)\})}\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(C)\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}_{0}(\underline{a}))\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(\mathcal{D}_{N}^{S})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}_{0}(\underline{a}))\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$